Parth (parthkohli):

A gas undergoes a process such that the ratio of P and V is constant and equal to one. What's the molar heat capacity?

Parth (parthkohli):

\[\]

Parth (parthkohli):

\[C_{m} = \frac{\partial q }{\partial T}\]Is this relation right?

OpenStudy (hwyl):

MATH PROCESSING ERROR

Parth (parthkohli):

That drawing isn't important. Just a PV graph with slope one.

Parth (parthkohli):

\[q = \Delta U - w\]\[= \frac{3}2RT + pV = \frac{3}{2}RT + RT =\frac{5}{2}RT\]

Parth (parthkohli):

I don't think it's 5R/2 though.

Parth (parthkohli):

Looks like I nowhere used the fact that \(p=V\).

OpenStudy (thomas5267):

I suspect \(w=\int p(t)V(t)\,dt=\int p(t)^2\,dt\).

Parth (parthkohli):

Oh, what a stupid mistake.

Parth (parthkohli):

\[\frac{pV}{2}=\frac{p^2}{2}= \frac{V^2}{2}\]

Parth (parthkohli):

well duh

Parth (parthkohli):

then we're done here.

OpenStudy (thomas5267):

Why?

OpenStudy (hwyl):

WHY IS MATH NOT PROCESSING

OpenStudy (hwyl):

\(TEST\)

OpenStudy (hwyl):

WTF

Parth (parthkohli):

just refresh

OpenStudy (hwyl):

K NOW IT IS BACK

OpenStudy (thomas5267):

So the answer is 3/2R?

Parth (parthkohli):

No. 2R.